# On the Lavrentiev gap for manifold-valued maps

### v1

@@ Abstract.
We investigate the validity and the failure of modular density of smooth maps on compact manifolds.
@@ 1. Introduction
(1.1) $\mathsf{M}\text{ be an $\textnormal{{m}}$-dimensional oriented compact manifold},$
(1.2) $\mathsf{N}\text{ be an $\textnormal{{n}}$-dimensional oriented compact manifold},$
and consider a Sobolev space $W^{1,\varphi}$. The problem of approximating Sobolev maps between manifolds is classical, and its resolution depends critically on both the function space and the topology of the manifolds. For maps with values in $\mathbb{R}^{N}$ and Young function $\varphi(t)=t^{p}$, the approximation argument is straightforward: standard convolution with a smooth kernel produces a sequence of smooth maps that converge strongly. When the target is a manifold $\mathsf{N}$, the situation is more delicate. Convolution now takes values in the convex hull of $\mathsf{N}$, so one must subsequently project these values back onto $\mathsf{N}$, for instance by using the nearest-point projection. This procedure works in the classical Sobolev setting $\varphi(t)=t^{p}$ for the superdimensional case $p\geq\textnormal{{m}}$. Indeed, by Morrey-Sobolev embedding, maps in $W^{1,p}(\mathsf{M},\mathsf{N})$ are continuous when $p>\textnormal{{m}}$ or belong to the space of functions with vanishing mean oscillation when $p=\textnormal{{m}}$ (see [8]), so projecting them back onto $\mathsf{N}$ then yields a sequence in $C^{\infty}(\mathsf{M},\mathsf{N})$ converging strongly in $W^{1,p}(\mathsf{M},\mathsf{N})$. The case $p<\textnormal{{m}}$ is more subtle, and the possibility of strong approximation depends on the topology of both manifolds. Even for the unit sphere $\textnormal{{S}}^{2}$ the problem is nontrivial. A classical example due to Schoen & Uhlenbeck [47] is the map
@@
can become unbounded on any ball $\textnormal{{B}}\subset\Omega$ intersecting the transition region $\{a(\cdot)=0\}$. Then, condition (1.10), is precisely designed to ensure that $a(\cdot)$ decays sufficiently rapidly to prevent uncontrolled degeneracy, thereby guaranteeing maximal gradient regularity for local minimizers. The necessity of such a condition has been confirmed through explicit counterexamples in [2, 4, 25, 26, 3] showing that, when (1.10) is violated, the so-called Lavrentiev phenomenon can occur, i.e., the infimum of a variational integral over smooth functions is strictly greater than its infimum over the natural energy space, and this prevents the density of smooth function in $W^{1,\varphi}$. We remark that, in our manifold-valued framework, the compactness of $\mathsf{N}$ ensures the boundedness of maps, which allows us to take $q\leq p+\alpha$, in the case $p\leq\textnormal{{m}}$.
@@
(1.16) $\textnormal{{m}}\geq 3\text{ and }\int^{\infty}\left(\frac{t}{\varphi^{-}_{\mathsf{M}}(t)}\right)^{\frac{2}{\textnormal{{m}}-2}}\left(\int_{t}^{\infty}\left(\frac{s}{\varphi^{-}_{\mathsf{M}}(s)}\right)^{\frac{1}{\textnormal{{m}}-2}}\right)^{-\textnormal{{m}}}\,{\rm d}t=\infty.$
@@
with $p\geq\textnormal{{m}}$, $0\leq a(\cdot)\in C^{0,\alpha}$, and $q$ satisfying (1.10). Variable exponent functionals of the type (1.11) are admitted, as long as $\textnormal{{m}}\leq p(x)\leq q(x)\leq q$, $0\leq a(\cdot)\in C^{0,\alpha}$, $p(\cdot),q(\cdot)$ are log-Hölder continuous and (1.10) is satisfied.
@@
Before proceeding, we make few observations regarding Theorem 1.2, emphasizing its connection with the classical Sobolev case $W^{1,p}$ and the role of topological assumptions.
@@
We observe that the situation of Theorem 1.2 reflects the one we have for the classical case $\varphi=t^{p}$ where, for $p<\textnormal{{m}}$, some topological assumptions on $\mathsf{M}$ and $\mathsf{N}$ are necessary for the density result, as shown, for instance, in [7, 30, 9, 34, 33, 35, 15]. Indeed, in Theorem 1.2, $W^{1,\varphi}$ can be a function space arbitrarily larger than $W^{1,\textnormal{{m}}}$, so Theorem 1.2 complements Theorem 1.1.
@@
$\inf_{W^{1,\varphi}(\mathsf{M},\mathsf{N})}\int_{\mathsf{M}}\varphi(x,|Du|)\,{\rm d}x=\inf_{C^{\infty}(\mathsf{M},\mathsf{N})}\int_{\mathsf{M}}\varphi(x,|Du|)\,{\rm d}x.$
@@
This condition is sharp for both $p\leq\textnormal{{m}}$ and $p>\textnormal{{m}}$ as highlighted in [2, Remark 35]. Following the fractal constructions of Balci, Diening & Surnachev [2], we show that when (1.19) holds, Lavrentiev phenomenon occurs and consequently smooth maps are not dense in $W^{1,\varphi}(\mathsf{M},\mathsf{N})$, for $\mathsf{M}:=[-1,1]^{\textnormal{{m}}}$, i.e. the m-dimensional cube, and $\mathsf{N}=\textnormal{{S}}^{N-1}_{{\Lambda}}$, where $\Lambda$ represents a suitable radii of the sphere; note that the sphere $\textnormal{{S}}^{N-1}_{{\Lambda}}$ is (N2)(N-2)-connected. This provides the first vectorial counterexample of the Lavrentiev phenomenon, which we further extend to manifold-valued maps. This is the content of the next theorem. First, let us introduce some notation. We set $Q_{\mathsf{M}}=(-1,1)^{\textnormal{{m}}}$, and for a given function $u_{0}\in C^{\infty}(\overline{Q}_{\mathsf{M}},\textnormal{{S}}^{N-1}_{\Lambda})$, we define the spaces
@@
(1.22) $\inf_{u\in W^{1,\varphi}_{u_{0}}(Q_{\mathsf{M}},\textnormal{{S}}^{N-1}_{\Lambda})}\int_{Q_{\mathsf{M}}}\varphi(x,|Du|)\,{\rm d}\mathscr{H}<\inf_{u\in C^{\infty}_{u_{0}}(\overline{Q_{\mathsf{M}}},\textnormal{{S}}^{N-1}_{\Lambda})}\int_{Q_{\mathsf{M}}}\varphi(x,|Du|)\,{\rm d}\mathscr{H}$
and there is no sequence $\{u_{\ell}\}_{\ell}\subset C^{\infty}_{u_{0}}(\overline{Q_{\mathsf{M}}},\textnormal{{S}}^{N-1}_{{\Lambda}})$ such that $\int_{Q_{\mathsf{M}}}\varphi(x,|Du_{\ell}-D\bar{u}|)\,{\rm d}\mathscr{H}\to 0$.
@@ 2.2. Properties of Young functions
(2.13) $\textnormal{{m}}\geq 3\text{ and }\int^{\infty}\left(\frac{t}{\Psi(t)}\right)^{\frac{2}{\textnormal{{m}}-2}}\left(\int_{t}^{\infty}\left(\frac{s}{\Psi(s)}\right)^{\frac{1}{\textnormal{{m}}-2}}\right)^{-\textnormal{{m}}}\,{\rm d}t=\infty.$

### v2

@@ Abstract.
We characterize necessary and sufficient conditions ensuring approximation by smooth maps in the Sobolev space $W^{1,\varphi}$ between suitable compact manifolds, where the prototypical example of Young function $\varphi$ is the double-phase integrand.
@@ 1. Introduction
(1.1) $\mathsf{M}\text{ be an $\textnormal{{m}}$-dimensional oriented compact Riemannian manifold}$
[…]
and consider the Sobolev space $W^{1,\varphi}$, where $\varphi$ is a Young function. The problem of approximating Sobolev maps between manifolds is classical, and its resolution depends critically on both the function space and the topology of the manifolds. For maps with values in $\mathbb{R}^{N}$ and Young function $\varphi(t)=t^{p}$, the approximation argument is straightforward: standard convolution with a smooth kernel produces a sequence of smooth maps that converge strongly. When the target is a manifold $\mathsf{N}$, the situation is more delicate. Convolution now takes values in the convex hull of $\mathsf{N}$, so one must subsequently project these values back onto $\mathsf{N}$, for instance by using the nearest-point projection. This procedure works in the classical Sobolev setting $\varphi(t)=t^{p}$ for the superdimensional case $p\geq\textnormal{{m}}$. Indeed, by Morrey-Sobolev embedding, maps in $W^{1,p}(\mathsf{M},\mathsf{N})$ are continuous when $p>\textnormal{{m}}$ or belong to the space of functions with vanishing mean oscillation when $p=\textnormal{{m}}$ (see [8]), so projecting them back onto $\mathsf{N}$ then yields a sequence in $C^{\infty}(\mathsf{M},\mathsf{N})$ converging strongly in $W^{1,p}(\mathsf{M},\mathsf{N})$. The case $p<\textnormal{{m}}$ is more subtle, and the possibility of strong approximation depends on the topology of both manifolds. Even for the unit sphere $\textnormal{{S}}^{2}$ the problem is nontrivial. A classical example due to Schoen & Uhlenbeck [48] is the map
@@
can become unbounded on any ball $\textnormal{{B}}\subset\Omega$ intersecting the transition region $\{a(\cdot)=0\}$. Then, condition (1.10), is precisely designed to ensure that $a(\cdot)$ decays sufficiently rapidly to prevent uncontrolled degeneracy, thereby guaranteeing maximal gradient regularity for local minimizers. The necessity of such a condition has been confirmed through explicit counterexamples in [2, 4, 25, 26, 3] showing that, when (1.10) is violated, the so-called Lavrentiev phenomenon can occur, i.e., the infimum of a variational integral over smooth functions is strictly greater than its infimum over the natural energy space, and this prevents the density of smooth functions in $W^{1,\varphi}$. We remark that, in our manifold-valued framework, the compactness of $\mathsf{N}$ ensures the boundedness of maps, which allows us to take $q\leq p+\alpha$, in the case $p\leq\textnormal{{m}}$.
@@
(1.16) $\textnormal{{m}}\geq 3\text{ and }\int^{\infty}\left(\frac{t}{\varphi^{-}_{\mathsf{M}}(t)}\right)^{\frac{2}{\textnormal{{m}}-2}}\left(\int_{t}^{\infty}\left(\frac{s}{\varphi^{-}_{\mathsf{M}}(s)}\right)^{\frac{1}{\textnormal{{m}}-2}}\,{\rm d}s\right)^{-\textnormal{{m}}}\,{\rm d}t=\infty.$
@@
with $p\geq\textnormal{{m}}$, $0\leq a(\cdot)\in C^{0,\alpha}$, and $q$ satisfying (1.10). Variable exponent functionals of the type (1.11) are also admitted, as long as $\textnormal{{m}}\leq p(x)\leq q(x)\leq q$, $0\leq a(\cdot)\in C^{0,\alpha}$, $p(\cdot),q(\cdot)$ are log-Hölder continuous and (1.10) is satisfied.
@@
Before proceeding, we make a few observations regarding Theorem 1.2, emphasizing its connection with the classical Sobolev case $W^{1,p}$ and the role of topological assumptions.
@@
We observe that the situation of Theorem 1.2 reflects the one occurring in the classical case $\varphi=t^{p}$ where, for $p<\textnormal{{m}}$, some topological assumptions on $\mathsf{M}$ and $\mathsf{N}$ are necessary for the density result, as shown, for instance, in [7, 30, 9, 34, 33, 35, 15]. Indeed, in Theorem 1.2, $W^{1,\varphi}$ can be a function space arbitrarily larger than $W^{1,\textnormal{{m}}}$, so Theorem 1.2 complements Theorem 1.1.
@@
$\inf_{W^{1,\varphi}(\mathsf{M},\mathsf{N})}\int_{\mathsf{M}}\varphi(x,|Du|)\,{\rm d}\mathscr{H}^{\textnormal{{m}}}=\inf_{C^{\infty}(\mathsf{M},\mathsf{N})}\int_{\mathsf{M}}\varphi(x,|Du|)\,{\rm d}\mathscr{H}^{\textnormal{{m}}}.$
@@
This condition is sharp for both $p\leq\textnormal{{m}}$ and $p>\textnormal{{m}}$ as highlighted in [2, Remark 35]. Following the fractal constructions of Balci, Diening & Surnachev [2], we show that when (1.19) holds, Lavrentiev phenomenon occurs and consequently smooth maps are not dense in $W^{1,\varphi}(\mathsf{M},\mathsf{N})$, for $\mathsf{M}:=[-1,1]^{\textnormal{{m}}}$, i.e. the m-dimensional cube, and $\mathsf{N}=\textnormal{{S}}^{N-1}_{{\Lambda}}$, where $\Lambda$ represents a suitable radius of the sphere; note that the sphere $\textnormal{{S}}^{N-1}_{{\Lambda}}$ is (N2)(N-2)-connected. This provides the first vectorial counterexample of the Lavrentiev phenomenon, which we further extend to manifold-valued maps. This is the content of the next theorem. First, let us introduce some notation. We set $Q_{\mathsf{M}}=(-1,1)^{\textnormal{{m}}}$, and for a given function $u_{0}\in C^{\infty}(\overline{Q}_{\mathsf{M}},\textnormal{{S}}^{N-1}_{\Lambda})$, we define the spaces
@@
(1.22) $\inf_{u\in W^{1,\varphi}_{u_{0}}(Q_{\mathsf{M}},\textnormal{{S}}^{N-1}_{\Lambda})}\int_{Q_{\mathsf{M}}}\varphi(x,|Du|)\,{\rm d}x<\inf_{u\in C^{\infty}_{u_{0}}(\overline{Q}_{\mathsf{M}},\textnormal{{S}}^{N-1}_{\Lambda})}\int_{Q_{\mathsf{M}}}\varphi(x,|Du|)\,{\rm d}x$
and there is no sequence $\{u_{\ell}\}_{\ell}\subset C^{\infty}_{u_{0}}(\overline{Q}_{\mathsf{M}},\textnormal{{S}}^{N-1}_{{\Lambda}})$ such that $\int_{Q_{\mathsf{M}}}\varphi(x,|Du_{\ell}-D\bar{u}|)\,{\rm d}x\to 0$.
@@ 2.2. Properties of Young functions
(2.13) $\textnormal{{m}}\geq 3\text{ and }\int^{\infty}\left(\frac{t}{\Psi(t)}\right)^{\frac{2}{\textnormal{{m}}-2}}\left(\int_{t}^{\infty}\left(\frac{s}{\Psi(s)}\right)^{\frac{1}{\textnormal{{m}}-2}}\,{\rm d}s\right)^{-\textnormal{{m}}}\,{\rm d}t=\infty.$
@@ 2.3. Musielak-Orlicz spaces
2.4. The double of the manifold $\mathsf{M}$.
Let $\mathsf{M}$ be an oriented, compact Riemannian m-manifold with smooth boundary $\partial\mathsf{M}\neq\emptyset$. We can consider its double denoted by $\mathcal{D}(\mathsf{M})$ as follows: we consider two copies of $\mathsf{M}$, denoted by $\mathsf{M}_{1},\mathsf{M}_{2}$ respectively, and we attach them along the boundary, so that $\partial\mathsf{M}_{1}=\partial\mathsf{M}_{2}=\partial\mathsf{M}$. The precise construction can be found in [40, Theorem 9.29, Example 9.32].
In particular, there exist two diffeomorphisms $\mathcal{i}_{1}:\mathsf{M}\to\mathsf{M}_{1}$, $\mathcal{i_{2}}:\mathsf{M}\to\mathsf{M}_{2}$ (often called identification maps between $\mathsf{M}$ and its $i$-th copy $\mathsf{M}_{i}$), and we have
(2.16) $\mathcal{D}(\mathsf{M})=\mathsf{M}_{1}\cup\mathsf{M}_{2},\quad\text{with}\quad\mathsf{M}_{1}\cap\mathsf{M}_{2}=\mathcal{i}_{1}(\partial\mathsf{M})=\partial\mathsf{M}_{1}=\mathcal{i}_{2}(\partial\mathsf{M})=\partial\mathsf{M}_{2}=\partial\mathsf{M}$
where the last equality is just formal and has to be intended in the sense of diffeomorphisms. This means that $\mathcal{D}(\mathsf{M})$ is made of the two copies of $\mathsf{M}$, whose interiors are disjoint, and such copies intersect only at the hypersurface $\partial\mathsf{M}$. Hence, by identifying $\mathsf{M}$ with one of its copies, say $\mathsf{M}_{1}$, we may effectively consider $\mathsf{M}\subset\mathcal{D}(\mathsf{M})$ as a submanifold.
Moreover, if $\mathsf{M}$ is a compact, oriented Riemannian m-dimensional manifold, so is the double $\mathcal{D}(\mathsf{M})$. The key feature is that
(2.17) $\partial\mathcal{D}(\mathsf{M})=\emptyset,\quad\text{and thus}\quad\partial\mathsf{M}\subset\mathcal{D}(\mathsf{M})=\mathrm{Int}\,\mathcal{D}(\mathsf{M})$
where $\mathrm{Int}$ denotes the interior of the manifold. Therefore, we have that $\partial\mathsf{M}$ is a $(\textnormal{{m}}-1)$-dimensional submanifold embedded into $\mathcal{D}(\mathsf{M})$.
Let us now consider a $\mathscr{H}^{\textnormal{{m}}}$-measurable function $v:\mathsf{M}\to\mathbb{R}^{N}$, $N\in\mathbb{N}$; its extension to the double $\mathcal{D}(\mathsf{M})$ is immediately defined by setting
(2.18) $\tilde{v}:\mathcal{D}(\mathsf{M})\to\mathbb{R}^{N},\quad\tilde{v}|_{\mathsf{M}_{1}}:=v\circ\mathcal{i}_{1},\quad\tilde{v}|_{\mathsf{M}_{2}}:=v\circ\mathcal{i}_{2}.$
so that $\tilde{v}|_{\mathsf{M}}=v$, modulo identifying $\mathsf{M}$ with $\mathsf{M}_{i}$. It immediately follows by definition that the images coindice $Im(v~)\mathrm$$\mathrm{Im}(\tilde{v})$, and if $v$ is real-valued
(2.19) $\sup_{\mathsf{M}}v=\sup_{\mathcal{D}(\mathsf{M})}\tilde{v},\quad\inf_{\mathsf{M}}v=\inf_{\mathcal{D}(\mathsf{M})}\tilde{v}\quad\text{and hence}\quad\operatorname*{osc}_{\mathsf{M}}v=\operatorname*{osc}_{\mathcal{D}(\mathsf{M})}\tilde{v}$.
Moreover, if $v\in W^{1,1}(\mathsf{M},\mathbb{R}^{N})$, then its extension $\tilde{v}\in W^{1,1}(\mathcal{D}(\mathsf{M}),\mathbb{R}^{N})$. To see this, first recall the following basic fact from the theory of Sobolev spaces: given two functions $w_{1}\in W^{1,1}(\Omega_{1}),w_{2}\in W^{1,1}(\Omega_{2})$ on two regular domains sets $\Omega_{1},\Omega_{2}\subset\mathbb{R}^{\textnormal{{m}}}$, such that $\Omega_{1}\cap\Omega_{2}=\Gamma$ is a regular hypersurface, then the function
$w(x)=\begin{cases}w_{1}(x)\quad x\in\Omega_{1}\\ w_{2}(x)\quad x\in\Omega_{2}\end{cases}$
belongs to the Sobolev spaces $W^{1,1}(\Omega_{1}\cup\Omega_{2})$ if and only if the traces $w_{1}|_{\Gamma}=w_{2}|_{\Gamma}$ coincide.
Hence, if $v\in W^{1,1}(\mathsf{M},\mathbb{R}^{N})$, then by (2.18) we have $v_{i}:=\tilde{v}|_{\mathsf{M}_{i}}\in W^{1,1}(\mathsf{M}_{i},\mathbb{R}^{N})$ for each $i=1,2$, and $v_{1}=v_{2}$ on $\mathsf{M}_{1}\cap\mathsf{M}_{2}=\partial\mathsf{M}$, so by (2.17) and from the preceding discussion (applied on local charts of $\mathcal{D}(\mathsf{M})$) we have $\tilde{v}\in W^{1,1}(\mathcal{D}(\mathsf{M}),\mathbb{R}^{N})$.
Now, let $\varphi:\mathsf{M}\times[0,\infty)\to[0,\infty)$ be a Young function as in (1.3)-(1.8), and define its extension to the double
[…]
It immediately follows from (2.18)-(2.19) that $\tilde{\varphi}$ still satisfies (1.3)-(1.8), and
(2.21) $\tilde{\varphi}_{\mathcal{D}(\mathsf{M})}^{-}(t)=\inf_{x\in\mathcal{D}(\mathsf{M})}\varphi(x,t)=\varphi^{-}_{\mathsf{M}}(t)\quad\text{for every $t\geq 0$.}$
Next, let $v\in W^{1,\varphi}(\mathsf{M},\mathbb{R}^{N})$, and let $\tilde{v}$ and $\tilde{\varphi}$ be as in (2.18)-(2.20). We claim that $\tilde{v}\in W^{1,\tilde{\varphi}}(\mathcal{D}(\mathsf{M}),\mathbb{R}^{N})$. Indeed, we saw that $\tilde{v}\in W^{1,1}(\mathcal{D}(\mathsf{M}),\mathbb{R}^{N})$, and using that $\mathsf{M}=\mathsf{M}_{j}$ (modulo diffeomorphism $\mathcal{i}_{j}$, $j=1,2$), together with (2.6) and (2.16), we deduce
$\int_{\mathcal{D}(\mathsf{M})}\tilde{\varphi}(x,|D\tilde{v}|)\,{\rm d}\mathscr{H}^{\textnormal{{m}}}=\int_{\mathsf{M}_{1}}\tilde{\varphi}(x,|D\tilde{v}|)\,{\rm d}\mathscr{H}^{\textnormal{{m}}}+\int_{\mathsf{M}_{2}}\tilde{\varphi}(x,|D\tilde{v}|)\,{\rm d}\mathscr{H}^{\textnormal{{m}}}\leq C\,\int_{\mathsf{M}}\varphi(x,|Dv|)\,{\rm d}\mathscr{H}^{\textnormal{{m}}}<\infty$
which shows that $\tilde{v}\in W^{1,\varphi}(\mathcal{D}(\mathsf{M}),\mathsf{N})$, and
$\int_{\mathcal{D}(\mathsf{M})}\tilde{\varphi}(x,|D\tilde{v}|)\,{\rm d}\mathscr{H}^{\textnormal{{m}}}\geq\int_{\mathsf{M}_{1}}\tilde{\varphi}(x,|D\tilde{v}|)\,{\rm d}\mathscr{H}^{\textnormal{{m}}}\geq c\,\int_{\mathsf{M}}\varphi(x,|Dv|)\,{\rm d}\mathscr{H}^{\textnormal{{m}}}.$
We sum up what we have just found in the following
Lemma 2.2 (Sobolev extension to the double $\mathcal{D}(\mathsf{M})$).
Let $\mathsf{M}$ be an oriented, compact m-dimensional Riemannian manifold, with smooth boundary $\partial\mathsf{M}\neq\emptyset$, and let $\varphi$ be a Young function fulfilling (1.3)-(1.8). Then, if $v\in W^{1,\varphi}(\mathsf{M},\mathbb{R}^{N})$, its extension $\tilde{v}$ to the double given by (2.18) belongs to $W^{1,\tilde{\varphi}}(\mathcal{D}(\mathsf{M}),\mathrm{Im}(v))$, with
(2.22) $c\,\|v\|_{W^{1,\varphi}(\mathsf{M},\mathbb{R}^{N})}\leq\|\tilde{v}\|_{W^{1,\tilde{\varphi}}(\mathcal{D}(\mathsf{M}),\mathbb{R}^{N})}\leq C\|v\|_{W^{1,\varphi}(\mathsf{M},\mathbb{R}^{N})},$
for some constants $c,C=c,C(\mathsf{M},N,\beta,\gamma)>0$, where $\tilde{\varphi}$ is given by (2.20) and satisfies the structural assumptions (1.3)-(1.8).
Remark 2.1.
In view of Lemma 2.2 and (2.21), it suffices to prove Theorems 1.1-1.2 in the case $\partial\mathsf{M}=\emptyset$. Indeed, assume that both theorems hold for $\varphi$ and when the reference manifold $\mathsf{M}$ has no boundary.
Then, if $\partial\mathsf{M}\neq\emptyset$, let $v\in W^{1,\varphi}(\mathsf{M},\mathsf{N})$ and $\tilde{v}\in W^{1,\tilde{\varphi}}(\mathcal{D}(\mathsf{M}),\mathsf{N})$ be its extension of Lemma 2.2. Since $\tilde{\varphi}$ still satisfies the assumptions of Theorems 1.1 (by (2.21)) and 1.2, $\partial\mathcal{D}(\mathsf{M})=\emptyset$, $\mathcal{D}(\mathsf{M})$ is a m-dimensional oriented compact Riemannian manifold, then by applying Theorems 1.1-1.2 in the boundary-less case, we may find a sequence $v_{\ell}\in C^{\infty}(\mathcal{D}(\mathsf{M}),\mathsf{N})$ such that $v_{\ell}\to\tilde{v}$ in $W^{1,\tilde{\varphi}}(\mathcal{D}(\mathsf{M}),\mathsf{N})$. Therefore, $v_{\ell}|_{\mathsf{M}}\xrightarrow{\ell\to\infty}v$ in $W^{1,\varphi}(\mathsf{M},\mathsf{N})$ owing to the identification $\mathsf{M}=\mathsf{M}_{1}\subset\mathcal{D}(\mathsf{M})$.
2.5. Vanishing web oscillations